Tim is twice as old as Tom. Last year Tim was 15 years older than Tom. WHat are their present ages?
Found 3 solutions by richwmiller, mananth, stanbon:
Answer by richwmiller(17219)

repeated or similar problem

Answer by mananth(16949)

let tom be x years
tim is 2x
last year
tom x-1
tim 2x-1
tim = 15+tom
2x-1= x-1+15
x=15 toms age
tim's age =30 years



Answer by stanbon(75887)

Tim is twice as old as Tom. Last year Tim was 15 years older than Tom. WHat are their present ages?
---
Equations:
Let Tim's age be "x"; Let Tom's age be "y":
x = 2y
(x-1)= (y-1) + 15
---
Substitute for "x" and solve for "y":
2y-1 = y+14
y = 15 (Tom's age now)
---
Since x = 2y, x = 30 (Tim's age now)
========================================
Cheers,
Stan H.
==============